More Info 2014 Jun 10 Adam Buckner and Amber Kwan have agreed to pool their assets and form a partnership to be called B&K Consuting. They agree to share all profits equally and make the following initial investments Cash Accounts receivable (net) Office furniture Buckner $ 15,000 33,000 36,000 Kwan 30,000 27,000 24,000 Dec 31 The partnerships reported net income was $195,000 for the year ended December 31, 2014 2015 Jan 1 Buckner and Kwan agree to accept Heidi Nguen into the partnership with a $180,000 investment for 30 percent of the business. The partnership agreement is amended to provide for the following sharing of profits and losses: Buckner 90,000 5% 2016 Annual salary Interest on capital balance Balance in ratio of The partnerships reported net income was $480,000. 120,000 75,000 5% Oct 10 Dec Jan 2 Buckner withdrew $84,000 cash from the partnership and Kwan withdrew $57,000. (Nguen did not make any withdrawals.) 5% artnerships reported net income was $255,000 Dec 31 After a disagreement as to the direction in which the partnership should be moving, Nguen decided to withdraw from the partnership. The three partners agreed that Nguen could take cash or 5300,000 in exchange for her equity in the partnership PrintDoneNeed help with 2016 Dec 31 journal entry.
Answer #1

Journal entries for December 31, 2016:

Date

Accounts Title and explanation

Debit($)

Credit($)

December 31,2016

Profit and Loss Account

255000

      Buckner’s Capital Account

76500

       Kwan’ s Capital Account

51000

      Nguen’ s Capital Account

127500

(Partnership reported profit distributed between all partner in their profit sharing ratio i.e. 3:2:5)

Workings:

Sharing of profit to Buckner’s = 255000*3/10 = 76500

Sharing of profit to Kwan’s = 255000*2/10 = 51000

Sharing of profit to Nguen’s = 255000*5/10 = 127500